OpenStudy (anonymous):

solve the system of equations and give the value of x... 7x+4y=-13.... -7x-11y=48.... ANSWER: x=-1?x=7?....x=-7?....x=1?

OpenStudy (anonymous):

7x+4y=-13 CAN u express y in terms of x from here?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

7x+4y=-13 4y=-13-7x y=(-13-7x)/4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

If you don't know how to isolate a variable, then you need to learn how to do it.

OpenStudy (anonymous):

Regardless, this system can be solved easily with elimination. Add the equations to eachother.

OpenStudy (anonymous):

But we do not need y

OpenStudy (anonymous):

The 7x and -7x should cancel out nicely, and you can solve for y then.

OpenStudy (anonymous):

ALTHOUGH WE ARE TRYING TO SOLVE FOR X.. CAN WE CANCEL X? SINCE THEY ARE OPPOSITES?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then solve for y and substitute it any o equation and solve for x to find x

OpenStudy (anonymous):

one*

OpenStudy (anonymous):

Y=5?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

SO X=1?

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

YES.... I THINK I COULDNT FIGURE IT OUT BECAUSE I LEFT OUT THE (-)...

OpenStudy (anonymous):

Now, find x

OpenStudy (anonymous):

RIGHT.. BUT AFTER ITS ALL FACTORED IN.. THE ANSWER FOR X IS 1?